Matt is one year older than five times Kelly's age. In three years, Matt will be only three times Kelly's age. How old are they now?
Answer by rm29924(97)

Let x = Kelly's current age
Matt's current age is 5x+1
Kelly's age in 3 years = x+3
Matt's age in 3 years = 5x+1+3
Matt's age in 3 years = 3 times Kelly's age in 3 years
5x+1+3 = 3(x+3)
5x+4=3x+9
subtract 4 from both sides
5x=3x+5
subtract 3x from both sides
2x=5
divide both sides by 2 to get x by itself
x=2.5
answer: Kelly is 2.5 years old
Matt is 5(2.5)+1 = 13.5 years old